Sol Summary – May 15th

Crew 281 Sol Summary Report 15-05-2023
Sol: 1
Summary Title: First EVA and Training
Author’s name: Ritupriya Patil (XO) and Megan Kane (Commander)
Mission Status: Nominal
Sol Activity Summary: We went out for our first EVAs where the crew of 5 was split in two EVA groups 3 members in each. The prep for this included doing a quick pre brief, health and water check done by Rachel Jones (HSO). We donned the radio equipment and the spacesuit.
EVA 1: Ritupriya Patil (XO, EVA Lead), Ana Pires (Crew Scientist), KC Shasteen (Crew Engineer).
EVA 2: Megan Kane (Commander, EVA Lead), Ana Pires (Crew Scientist), Rachel Jones (HSO).
EVA prep began at 7:45 am when we started prepping our suits, radio comms, helmets and air circulation. KC SHashteen (Crew Engineer) made the pre-EVA checks with radios and spacesuit. We got suited and followed a 5 minute depressurization in the air lock, starting out at 8:30 am. We took the Perseverance and Curiosity rover to Marble Ritual for our first EVA and did a 100 m of training walk with the spacesuit and geotechnical equipment. We performed communication checks with the hab com and the EVA com. The crew successfully achieved another training goal of the EVA by recording 200 measurements for rock hardness with Schmidt Hammer and Equotip. In order to acquire images of the first scanline and take pictures of the geological features the crew scientist left the fibreglass tape onsite. This was a lesson learned and after the debrief, it was decided that no equipment from the EVA should be left behind. This EVA enabled us to work on design suggestions for future traverse planning to carry out the geological-geotechnical characterization. The crew returned ahead of time at 10:00 am.
The second EVA was performed after a 30 minute break from the first one since we had one crew scientist going back for the second EVA. The second crew started EVA at 10:51 am also following the depressurization in the airlock.Crew of 3 drove Spirit and Opportunity to Cow Dung road and stopped at Marble Ritual. Some tourists arrived at the EVA location and the crew lost some EVA time since they had to wait. Due to this, we truncated the EVA and were only able to do visual observations of the geological features and we were not able to document those.
In the afternoon, the Cacao (Chocolate) sprouting experiment was started in the Science Dome and the passive watering system was installed in the greenhab. Also in the afternoon cleaning fo the lower deck and repairs to the tunnel system were done.
Look Ahead Plan:
We have a plan for scouting EVA around the hab tomorrow which was approved
Additional experiments are in the works and existing experiments will be continued.
Anomalies in work: None
Weather: Fine
Crew Physical Status: Fine
EVA: Completed 2 today
Reports to be filed:
Greenhab Report
EVA Reports EVAs 1 and 2
Ops Report
Commander’s Report
Support Requested:
Support with Wifi password was requested and received
EVA assistance was requested and received

Copyright © The Mars Society. All rights reserved. | Main Site